I need help writing an equation in this problem please.
A car rental company's standard charge includes an initial fee plus an additional fee for each mile driven. The standard charge S (in dollars) is given by the function S= 14.95+0.60M, where M is the number of miles driven.

The company also offers an option to insure the car against damage. The insurance charge I (in dollars) is given by the function I= 5.70+0.15M.
Let C be the total charge (in dollars) for a rental that includes insurance. Write an equation relating C to M. Simplify your answer as much as possible.

Found 2 solutions by ikleyn, Theo:
Answer by ikleyn(52781)

.

In this problem, the total charge is the sum of two partial charges  C = S + I,  or


    C = (14.95 + 0.6M) + (5.70 + 0.15M) = 20.65 + 0.75M.



It is the equation you are asked for.

Solved, answered and carefully explained.



Answer by Theo(13342)

standard charge is S = 14.95 + .6M
insurance charge is I = 5.70 + .15M

C = total charge that includes standard charge plus insurance charge.

C = S + I = 14.95 + .6M + 5.70 + .15M

combine like terms to get:

C = 20.65 + .75M

for example, assume 100 miles, then:

S = 14.95 + 60 = 74.95
I = 5.70 + 15 = 20.70

S + I = 74.95 + 20.70 = 95.65

C = 20.65 + 75 = 95.65

that's your solution.